# 2cm2cm1cm1cm

### v1

@@ 4.5.1 Unbound pair scattering channel
is a trace-class operator, where we recall that $P_{x}$ is the orthogonal projection on the one-dimensional subspace $\mathbb{C}\mathfrak{\hat{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. By [34, Theorem 4.4, Chapter X], it follows in this case that
@@
Proof. Note that $P_{\mathrm{ac}}(\mathrm{H}_{f})=\mathfrak{1}$, thanks to Lemma 4.20. By (49), it justifies the strong limit given in the lemma. As we discuss in the proof of Proposition 4.1, for any $k\in\mathbb{T}^{2}$, the operator $A(k)$ is the sum of $M_{\mathfrak{f}\left(k\right)}\oplus A_{2,2}(k)$ and a compact operator $T$. In fact, as the function $\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ (defining the fiber Hamiltonian $A(k)$) is, absolutely summable (see (7)), the operator difference $T$ is even trace-class. As explained in Lemma 4.20, $R\left(\mathrm{V},\mathrm{v}\right)$ is also a trace-class operator, because $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ is absolutely summable, again by assumption. By using the Kato-Rosenblum theorem [35, Theorem XI.8], it thus follows that the wave operators
@@
$W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)=\left({\int_{\mathbb{T}^{2}}^{\oplus}}W^{\pm}\left(A\left(k\right),H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}$
@@
$\mathrm{ran}\left(W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\right)=\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\,\nu\left(\mathrm{d}k\right)\ .$
@@
We can then apply Proposition 5.6 together with Lemmata 4.20, 4.21, Theorem 5.3 and Equations (129)–(131) to arrive at
$W^{\pm}\left(UHU^{\ast},M_{\mathfrak{f}};\mathfrak{U}\right)$ $\doteq s-\lim_{t\rightarrow\mp\infty}\mathrm{e}^{itUHU^{\ast}}\mathfrak{U}\mathrm{e}^{-it\mathrm{H}_{f}}P_{\mathrm{ac}}\left(\mathrm{H}_{f}\right)$
$=s-\lim_{t\rightarrow\mp\infty}\mathrm{e}^{itUHU^{\ast}}\mathfrak{U}\mathrm{e}^{-it\mathrm{H}_{f}}$
$=s-\lim_{t\rightarrow\mp\infty}\mathrm{e}^{itUHU^{\ast}}\mathrm{e}^{-itH^{(1)}}\mathfrak{U}$
@@
$\mathrm{ran}\left(W^{\pm}\left(UHU^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\right)=\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\,\nu\left(\mathrm{d}k\right)\ .$
@@
as a Dyson series for all bounded operators $X,Y$. This can be applied to $X=H^{(1)}$ and $Y=UHU^{\ast}-X$, or in each fiber $k\in\mathbb{T}^{2}$ to $X=H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)$ and $Y=A\left(k\right)-X$. When $\mathrm{U}=\mathrm{V}\in\mathbb{R}_{0}^{+}$ and $\mathrm{v}=\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ in $H^{(1)}$, this result is particularly advantageous because the operator family $(Y_{t})_{t\in\mathbb{R}}$ appearing in Lemma 5.1 can be represented in a relatively simple way in this situation:
@@
$\mathrm{e}^{itH^{(1)}\left(\mathrm{U},\mathrm{u}\right)}\mathrm{e}^{i\left(s-t\right)UHU^{\ast}}\mathrm{e}^{-isH^{(1)}\left(\mathrm{U},\mathrm{u}\right)}\\ =\int_{\mathbb{T}^{2}}^{\oplus}\left(\begin{array}[]{cc}\cos_{\succ}\left(B_{1,2}\left(k\right)B_{2,1}\left(k\right);s,t\right)&-i\sin_{\succ}\left(B_{1,2}\left(k\right)B_{2,1}\left(k\right);s,t\right)\\ -i\sin_{\succ}\left(B_{2,1}\left(k\right)B_{1,2}\left(k\right);s,t\right)&\cos_{\succ}\left(B_{2,1}\left(k\right)B_{1,2}\left(k\right);s,t\right)\end{array}\right)\nu\left(\mathrm{d}k\right)\ ,$
@@
Proof. Let $H^{(1)}\equiv H^{(1)}\left(\mathrm{U},\mathrm{u}\right)$. We infer from Lemma 5.1 applied to $X=H^{(1)}$ and $Y=UHU^{\ast}-X$ that
$\mathrm{e}^{itH^{(1)}}\mathrm{e}^{i\left(s-t\right)UHU^{\ast}}\mathrm{e}^{-isH^{(1)}}=V_{t,s}\doteq\mathbf{1}+\sum_{n=1}^{\infty}\left(-i\right)^{n}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{n-1}}\mathrm{d}\tau_{n}B^{(\tau_{1})}\cdots B^{(\tau_{n})}\ ,$
@@
$B^{(t)}$ $=$ $\mathrm{e}^{itH^{(1)}}\left(UHU^{\ast}-H^{(1)}\right)\mathrm{e}^{-itH^{(1)}}$
@@ 4.5.2 Bound pair scattering channel
We start by studying the wave operator (48) with respect to the operators $X=UHU^{\ast}$ and $Y=M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$ (60), the identification operator $J$ being $\mathfrak{P}_{\mathrm{U}}$ (59) for any fixed Hubbard coupling constant $\mathrm{U}\in\mathbb{R}_{0}^{+}$.
@@
Let $h_{b}\in[0,1/2]$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$. Then, $UHU^{\ast}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$ and for every bounded continuous function $f\in C_{b}(\mathbb{R})$,
$\,f\left(UHU^{\ast}\right)\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}{\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\mathrm{U},k\right)\right)\nu\left(\mathrm{d}k\right)\text{ }.$
@@
$\left(UHU^{\ast}\mathfrak{P}_{\mathrm{U}}\varphi\right)\left(k\right)=A\left(k\right)\left(\mathfrak{P}_{\mathrm{U}}\varphi\right)\left(k\right)=\mathrm{E}\left(\mathrm{U},k\right)\varphi\left(k\right)\left\|\Psi\left(\mathrm{U},k\right)\right\|^{-1}\Psi\left(\mathrm{U},k\right)=\left(\mathfrak{P}_{\mathrm{U}}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\varphi\right)\left(k\right)\ ,$
i.e., $UHU^{\ast}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$, keeping in mind Equations (59) and (60). We then obtain the last assertion by using the Stone-Weierstrass theorem and the spectral theorem.
@@
$s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,f\left(UHU^{\ast}\right)\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\infty}{\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\infty,k\right)\right)\nu\left(\mathrm{d}k\right)\text{ }.$
@@
$s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,f\left(UHU^{\ast}\right)\mathfrak{P}_{\mathrm{U}}$ $=$ $s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,\mathfrak{P}_{\mathrm{U}}f\left(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\right)$

### v2

@@ 4.6.1 Unbound pair scattering channel
is a trace-class operator, where we recall that $P_{x}$ is the orthogonal projection on the one-dimensional subspace $\mathbb{C}\mathfrak{\hat{e}}_{x}\subseteq L^{2}(\mathbb{T}^{2})$. By [40, Theorem 4.4, Chapter X], it follows in this case that
@@
Proof. Note that $P_{\mathrm{ac}}(\mathrm{H}_{f})=\mathfrak{1}$, thanks to Lemma 4.21. By (55), it justifies the strong limit given in the lemma. As we discuss in the proof of Proposition 4.2, for any $k\in\mathbb{T}^{2}$, the operator $A(k)$ is the sum of $M_{\mathfrak{f}\left(k\right)}\oplus A_{2,2}(k)$ and a compact operator $T$. In fact, as the function $\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}$ (defining the fiber Hamiltonian $A(k)$) is, absolutely summable (see (7)), the operator difference $T$ is even trace-class. As explained in Lemma 4.21, $R\left(\mathrm{V},\mathrm{v}\right)$ is also a trace-class operator, because $\mathrm{v}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ is absolutely summable, again by assumption. By using the Kato-Rosenblum theorem [41, Theorem XI.8], it thus follows that the wave operators
@@
$W^{\pm}\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)=\left({\int_{\mathbb{T}^{2}}^{\oplus}}W^{\pm}\left(A\left(k\right),H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)\right)\nu\left(\mathrm{d}k\right)\right)\mathfrak{U}$
@@
$\mathrm{ran}\left(W^{\pm}\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\right)=\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\,\nu\left(\mathrm{d}k\right)\ .$
@@
We can then apply Proposition 5.6 together with Lemmata 4.21, 4.22, Theorem 5.3 and Equations (145)–(147) to arrive at
$W^{\pm}\left(\mathbb{U}H\mathbb{U}^{\ast},M_{\mathfrak{f}};\mathfrak{U}\right)$ $\doteq s-\lim_{t\rightarrow\mp\infty}\mathrm{e}^{it\mathbb{U}H\mathbb{U}^{\ast}}\mathfrak{U}\mathrm{e}^{-it\mathrm{H}_{f}}P_{\mathrm{ac}}\left(\mathrm{H}_{f}\right)$
$=s-\lim_{t\rightarrow\mp\infty}\mathrm{e}^{it\mathbb{U}H\mathbb{U}^{\ast}}\mathfrak{U}\mathrm{e}^{-it\mathrm{H}_{f}}$
$=s-\lim_{t\rightarrow\mp\infty}\mathrm{e}^{it\mathbb{U}H\mathbb{U}^{\ast}}\mathrm{e}^{-itH^{(1)}}\mathfrak{U}$
@@
$\mathrm{ran}\left(W^{\pm}\left(\mathbb{U}H\mathbb{U}^{\ast},\mathrm{H}_{f};\mathfrak{U}\right)\right)=\int_{\mathbb{T}^{2}}^{\oplus}L^{2}\left(\mathbb{T}^{2}\right)\oplus\{0\}\,\nu\left(\mathrm{d}k\right)\ .$
@@
as a Dyson series for all bounded operators $X,Y$. This can be applied to $X=H^{(1)}$ and $Y=\mathbb{U}H\mathbb{U}^{\ast}-X$,or in each fiber $k\in\mathbb{T}^{2}$ to $X=H^{(1)}\left(k\right)\oplus A_{2,2}\left(k\right)$ and $Y=A\left(k\right)-X$. When $\mathrm{U}=\mathrm{V}\in\mathbb{R}_{0}^{+}$ and $\mathrm{v}=\mathrm{u}:\mathbb{Z}^{2}\rightarrow\mathbb{R}_{0}^{+}$ in $H^{(1)}$, this result is particularly advantageous because the operator family $(Y_{t})_{t\in\mathbb{R}}$ appearing in Lemma 5.1 can be represented in a relatively simple way in this situation:
@@
$\mathrm{e}^{itH^{(1)}\left(\mathrm{U},\mathrm{u}\right)}\mathrm{e}^{i\left(s-t\right)\mathbb{U}H\mathbb{U}^{\ast}}\mathrm{e}^{-isH^{(1)}\left(\mathrm{U},\mathrm{u}\right)}\\ =\int_{\mathbb{T}^{2}}^{\oplus}\left(\begin{array}[]{cc}\cos_{\succ}\left(B_{1,2}\left(k\right)B_{2,1}\left(k\right);s,t\right)&-i\sin_{\succ}\left(B_{1,2}\left(k\right)B_{2,1}\left(k\right);s,t\right)\\ -i\sin_{\succ}\left(B_{2,1}\left(k\right)B_{1,2}\left(k\right);s,t\right)&\cos_{\succ}\left(B_{2,1}\left(k\right)B_{1,2}\left(k\right);s,t\right)\end{array}\right)\nu\left(\mathrm{d}k\right)\ ,$
@@
Proof. Let $H^{(1)}\equiv H^{(1)}\left(\mathrm{U},\mathrm{u}\right)$. We infer from Lemma 5.1 applied to $X=H^{(1)}$ and $Y=\mathbb{U}H\mathbb{U}^{\ast}-X$ that
$\mathrm{e}^{itH^{(1)}}\mathrm{e}^{i\left(s-t\right)\mathbb{U}H\mathbb{U}^{\ast}}\mathrm{e}^{-isH^{(1)}}=V_{t,s}\doteq\mathbf{1}+\sum_{n=1}^{\infty}\left(-i\right)^{n}\int_{s}^{t}\mathrm{d}\tau_{1}\cdots\int_{s}^{\tau_{n-1}}\mathrm{d}\tau_{n}B^{(\tau_{1})}\cdots B^{(\tau_{n})}\ ,$
@@
$B^{(t)}$ $=$ $\mathrm{e}^{itH^{(1)}}\left(\mathbb{U}H\mathbb{U}^{\ast}-H^{(1)}\right)\mathrm{e}^{-itH^{(1)}}$
@@ 4.6.2 Bound pair scattering channel
We start by studying the wave operator (54) with respect to the operators $X=\mathbb{U}H\mathbb{U}^{\ast}$ and $Y=M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$ (70), the identification operator $J$ being $\mathfrak{P}_{\mathrm{U}}$ (69) for any fixed Hubbard coupling constant $\mathrm{U}\in\mathbb{R}_{0}^{+}$.
@@
Let $h_{b}\in[0,1/2]$ and $\mathrm{U}\in\mathbb{R}_{0}^{+}$. Then, $\mathbb{U}H\mathbb{U}^{\ast}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$ and for every bounded continuous function $f\in C_{b}(\mathbb{R})$,
$\,f\left(\mathbb{U}H\mathbb{U}^{\ast}\right)\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}{\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\mathrm{U},k\right)\right)\nu\left(\mathrm{d}k\right)\text{ }.$
@@
$\left(\mathbb{U}H\mathbb{U}^{\ast}\mathfrak{P}_{\mathrm{U}}\varphi\right)\left(k\right)=A\left(k\right)\left(\mathfrak{P}_{\mathrm{U}}\varphi\right)\left(k\right)=\mathrm{E}\left(\mathrm{U},k\right)\varphi\left(k\right)\left\|\Psi\left(\mathrm{U},k\right)\right\|^{-1}\Psi\left(\mathrm{U},k\right)=\left(\mathfrak{P}_{\mathrm{U}}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\varphi\right)\left(k\right)\ ,$
i.e., $\mathbb{U}H\mathbb{U}^{\ast}\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\mathrm{U}}M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}$,keeping in mind Equations (69) and (70). We then obtain the last assertion by using the Stone-Weierstrass theorem and the spectral theorem.
@@
$s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,f\left(\mathbb{U}H\mathbb{U}^{\ast}\right)\mathfrak{P}_{\mathrm{U}}=\mathfrak{P}_{\infty}{\int_{\mathbb{T}^{2}}^{\oplus}}f\left(\mathrm{E}\left(\infty,k\right)\right)\nu\left(\mathrm{d}k\right)\text{ }.$
@@
$s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,f\left(\mathbb{U}H\mathbb{U}^{\ast}\right)\mathfrak{P}_{\mathrm{U}}$ $=$ $s-{\lim\limits_{\mathrm{U}\rightarrow\infty}}\,\mathfrak{P}_{\mathrm{U}}f\left(M_{\mathrm{E}\left(\mathrm{U},\cdot\right)}\right)$